A cannon tilted at 25 above the horizontal is fired from the top of a cliff 100 m exits the cannon at a speed of 150 m/s. The ground below is flat. tall. The ball Wh at is the highest point above the ground the ball will reach? What is the speed of the ball at the highest point? How long does it take the ball to reach the ground? How far horizontally will the ball have traveled when it reaches the ground? What is the speed of the ball as it hits the ground? What is the direction of the balls velocity when it hits the ground? a. b. c. d. e. f.
Answer #1

Given θ= 250

Initial velocity, u-150m/s

Initial horizontal velocity,ucosố 150 * cos25

u-136m/s UT

Initial vertical velocity,u_{y}=usin heta =150*sin25

uy-63.4m/s

=====================

Consider the vertical motion ,Rising Part

Final vertical velocity =0m/s

Use formula 205

2l 2gy

(Om/s)2 - (63.4m/s)2--2 9.81m/s2 y 63.4m/s) 2 *

-63.42-_2 * 9.81 * y

y=204.87m

From Ground ,height H=100m+y=100m+204.87m =304.87m

a. ANSWER: H-304.87m

====================

at Highest point v_{y}=0m/s and v_{x}=u_{x}=136m/s (there is no acceleration in horizontal direction)

v_{H}=sqrt{v_{x}^{2}+v_{y}^{2}}=sqrt{0^{2}+v_{y}^{2}}=v_{y}

b.ANSWER:{color{Red} v_{H}=136m/s}

=====================

c.

Consider vertical motion,Rising Part

Use Formula v=u+at

v_{y}=u_{y}-gt_{1}

0m/s=63.4m/s-9.81m/s^{2}*t_{1}

{color{Red} t_{1}=6.46s}

=========

Falling Part

Use Formula sut +1/2at2

H=0m/s*t_{2}+1/2*g*t_{2}^{2}

304.87m=0+0.5*9.81*t_{2}^{2}

304.87m=4.905t_{2}^{2}

{color{Red} t_{2}=7.88s}

t=t_{1}+t_{2}=6.46s+7.88s

c.ANSWER: {color{Red} t=14.34s}

=====================

d.

Speed =Distance/Time

Distance=Speed*Time

Since there is no acceleration in horizontal direction ,horizontal velocity remains same

X=u_{x}*t

X=136m/s*14.34s

d.ANSWER: {color{Red} X=1950.24m}

==========================

e.

Consider vertical motion,Falling Part

Use Formula 205

v_{y}^{2}-u_{y}^{2}=2gH

v_{y}^{2}-(0m/s)^{2}=2*9.81m/s^{2}*304.87m

v_{y}=sqrt{2*9.81*304.87}

y 77.34m/s(downwards)

{color{Red} v_{x}=136m/s}

v=sqrt{v_{x}^{2}+v_{y}^{2}}=sqrt{(136m/s)^{2}+(77.34m/s)^{2}}

e.ANSWER: {color{Red} v=156.45m/s}

=========================

f.

heta =tan^{-1}(v_{y}/v_{x})

heta =tan^{-1}(77.34/136)

f.ANSWER: {color{Red} heta =29.63^{circ}} below horizontal axis

===========================================